# TiC-LM: A Web-Scale Benchmark for Time-Continual LLM Pretraining

### v2

@@ 7 Conclusion
This paper presents work whose goal is to advance the efficiency of training and updating large language models. There are many potential societal consequences of our work. These improvements could help democratize LM development by enabling resource-constrained organizations to participate in research and development, while reducing the environmental footprint of model training through decreased energy consumption. However, increased accessibility of LM training also increases any risks that pertain to language models as a technology, such as the potential for malicious actors to spread disinformation.
@@ Acknowledgements
We are highly grateful to the many developers of the OpenLM library which was used extensively in this work. In particular, we’d like to thank George Smyrnis, Achal Dave, and Samir Yitzhak Gadre for helpful discussions and suggestions regarding its usage. We’d also like to thank Chun-Liang Li, Pavan Kumar Anasosalu Vasu, Preetum Nakkiran, and Russ Webb for helpful comments on earlier versions of the manuscript.
@@ B.4 Evaluation metrics
For some downstream evaluations, the evaluation periods do not exactly align with our chosen set of model checkpoints, making the definition of ID more nuanced. For such tasks, we define $a_{i}$ as the index of the nearest evaluation timestep that comes before the training timestep $i$, counting $R_{i,a_{i}}$ towards the ID average only if no other training timestep is closer to $a_{i}$ (i.e., $a_{i}\neq a_{i-1}$). Otherwise, we count $R_{i,j}$ towards backward and forward transfer when $j<a_{i}$ and $j\geq a_{i}$ respectively. For such evaluations, we define $a_{i}$ as the index of the evaluation timestep nearest to the training timestep $i$. Then ID performance is defined as $\sum_{i=1}^{T_{t}}=R_{i,a_{i}}/T_{t}$, backward transfer as $\sum_{i=1}^{T_{t}}\sum_{j\in S_{i}}\frac{R_{i,j}}{T_{t}|S_{i}|}$ where $S_{i}=\{j<a_{i}\}$, and forward transfer is defined similarly using the set $\{j>a_{i}\}$.
@@ D.5 TiC-StackE-English
Table 8: Results for TiC-StackE-English (440B token scale)
@@
Cyclic Cosine 440B 0.033 0.046 0.072
Cyclic Cosine + AR 440B 0.029 0.044 0.072
Cyclic Rsqrt 440B 0.029 0.043 0.071
ScheduleFree 440B 0.031 0.045 0.073
Replay ($\alpha=1/t$) 440B 0.031 0.054 0.077
Replay ($\alpha=1/2$) 440B 0.031 0.045 0.075
Replay ($\alpha=1/t$) + AR 440B 0.031 0.052 0.078
Replay ($\alpha=1/2$) + AR 440B 0.027 0.044 0.073
LwF 440B 0.033 0.044 0.072
EWC 440B 0.031 0.044 0.072
Oracle Series 1.16T 0.021 0.037 0.074
D.6 Results for 220B scale
Table 9: Downstream evaluations at 220B token scale. For all dynamic evaluations, we report perplexity values relative to Oracle-2024-07 with log-scaling. Meanwhile, Core is an average of the accuracies of 22 downstream zero/few-shot tasks used by Li et al. (2024a), evaluated only on the final model checkpoint (without scaling by an Oracle). Bold values are within one standard deviation (estimated with 3 runs of Cyclic Cosine) of the best in each column. Highlighted values indicate matching the Oracle series (within one standard deviation).
[…]
Bwd. ID Fwd. Bwd. ID Fwd. Bwd. ID Fwd. Bwd. ID Fwd.
Cyclic Cosine (std) 220B 0.033 0.052 0.085 0.039 0.052 0.072 0.041 0.078 0.156 0.057 0.025 0.217
Cyclic Cosine + AR 220B 0.033 0.054 0.087 0.035 0.051 0.074 0.032 0.077 0.159 0.084 0.052 0.228
Cyclic Rsqrt 220B 0.031 0.051 0.084 0.035 0.050 0.070 0.034 0.076 0.158 0.062 0.029 0.220
Schedule-Free 220B 0.035 0.055 0.087 0.040 0.055 0.074 0.038 0.079 0.160 0.084 0.051 0.236
Replay ($\alpha_{t}=1/t$) 220B 0.038 0.063 0.091 0.035 0.056 0.074 0.075 0.121 0.191 0.175 0.138 0.275
Replay ($\alpha_{t}=1/2$) 220B 0.032 0.055 0.086 0.034 0.053 0.072 0.055 0.094 0.170 0.099 0.069 0.237
Replay ($\alpha_{t}=1/t$) + AR 220B 0.039 0.063 0.092 0.034 0.055 0.077 0.066 0.119 0.193 0.197 0.169 0.277
Replay ($\alpha_{t}=1/2$) + AR 220B 0.033 0.057 0.088 0.033 0.052 0.074 0.047 0.096 0.176 0.129 0.098 0.246
LwF 220B 0.058 0.028 0.214 0.039 0.053 0.072 0.037 0.075 0.155 0.058 0.028 0.214
EWC 220B 0.030 0.051 0.083 0.034 0.050 0.069 0.033 0.077 0.162 0.070 0.040 0.222
Oracle Series 1.16T 0.021 0.052 0.082 0.018 0.042 0.062 0.023 0.074 0.151 0.047 0.000 0.218
Bwd. ID Fwd. Bwd. ID Fwd. Core (DCLM)
Cyclic Cosine (std) 220B 0.036 0.023 0.014 0.073 0.096 0.072 48.5
Cyclic Cosine + AR 220B 0.011 0.006 0.003 0.054 0.074 0.062 48.5
Cyclic Rsqrt 220B 0.025 0.015 0.009 0.066 0.092 0.071 49.0
Schedule-Free 220B 0.035 0.023 0.017 0.069 0.100 0.080 48.8
Replay ($\alpha_{t}=1/t$) 220B -0.009 -0.010 -0.006 0.054 0.046 0.057 48.9
Replay ($\alpha_{t}=1/2$) 220B 0.010 0.003 0.001 0.058 0.066 0.060 49.0
Replay ($\alpha_{t}=1/t$) + AR 220B -0.019 -0.015 -0.008 0.040 0.045 0.057 49.0
Replay ($\alpha_{t}=1/2$) + AR 220B -0.001 -0.004 -0.002 0.034 0.050 0.052 49.2
LwF 220B 0.034 0.021 0.013 0.076 0.104 0.073 48.5
EWC 220B 0.016 0.009 0.006 0.055 0.081 0.067 48.9
Oracle Series 1.16T -0.024 -0.028 -0.021 0.012 0.013 0.017 50.6
D.7 Extended replay exploration
Table 10: Further exploration of replay variants. Results are for the 220B token scale.
@@ D.7 Extended replay exploration
Replay ($\alpha=1/t$) 0.038 0.063 0.091 0.035 0.056 0.074
Replay (Exp) 0.032 0.054 0.086 0.037 0.052 0.072
Replay ($\alpha=0.1$) 0.036 0.061 0.090 0.035 0.055 0.075
Replay ($\alpha=0.3$) 0.034 0.058 0.088 0.034 0.054 0.074
Replay ($\alpha=0.5$) 0.032 0.055 0.086 0.034 0.053 0.072
Replay ($\alpha=0.7$) 0.032 0.053 0.085 0.036 0.053 0.072
Replay ($\alpha=0.9$) 0.033 0.052 0.085 0.039 0.054 0.072
Replay ($\alpha=1/t$) + AR 0.039 0.063 0.092 0.034 0.055 0.077
Replay (Exp) + AR 0.033 0.056 0.088 0.034 0.052 0.073
Replay ($\alpha=0.1$) + AR 0.038 0.062 0.091 0.034 0.054 0.076
Replay ($\alpha=0.3$) + AR 0.036 0.059 0.090 0.033 0.053 0.075
Replay ($\alpha=0.5$) + AR 0.033 0.057 0.088 0.033 0.052 0.074
Replay ($\alpha=0.7$) + AR 0.033 0.055 0.087 0.033 0.051 0.073
Replay ($\alpha=0.9$) + AR 0.033 0.053 0.086 0.035 0.052 0.072
@@
Replay ($\alpha=1/t$) 0.075 0.121 0.191 -0.009 -0.010 0.006
Replay (Exp) 0.054 0.093 0.170 0.026 0.012 0.006
Replay ($\alpha=0.1$) 0.062 0.106 0.183 -0.007 -0.009 -0.004
Replay ($\alpha=0.3$) 0.060 0.103 0.176 -0.001 -0.002 -0.001
Replay ($\alpha=0.5$) 0.055 0.094 0.170 0.010 0.003 0.001
Replay ($\alpha=0.7$) 0.050 0.091 0.169 0.018 0.009 0.005
Replay ($\alpha=0.9$) 0.042 0.081 0.160 0.028 0.016 0.007
Replay ($\alpha=1/t$) + AR 0.066 0.119 0.193 -0.019 -0.015 -0.008
Replay (Exp) + AR 0.044 0.094 0.174 0.005 -0.001 -0.000
Replay ($\alpha=0.1$) + AR 0.057 0.110 0.188 -0.018 -0.014 -0.005
Replay ($\alpha=0.3$) + AR 0.052 0.103 0.179 -0.012 -0.009 -0.004
Replay ($\alpha=0.5$) + AR 0.047 0.096 0.176 -0.006 -0.006 -0.002
Replay ($\alpha=0.7$) + AR 0.042 0.091 0.170 -0.001 -0.003 -0.002
Replay ($\alpha=0.9$) + AR 0.038 0.085 0.165 0.005 $0$0.001
@@
Replay ($\alpha=1/t$) 0.054 0.046 0.057 0.175 0.138 0.298
Replay (Exp) 0.070 0.063 0.063 0.073 0.041 0.233
Replay ($\alpha=0.1$) 0.042 0.045 0.051 0.165 0.146 0.291
Replay ($\alpha=0.3$) 0.056 0.059 0.057 0.132 0.111 0.280
Replay ($\alpha=0.5$) 0.058 0.066 0.060 0.099 0.069 0.271
Replay ($\alpha=0.7$) 0.066 0.083 0.068 0.090 0.055 0.266
Replay ($\alpha=0.9$) 0.069 0.078 0.064 0.070 0.035 0.264
Replay ($\alpha=1/t$) + AR 0.040 0.045 0.057 0.197 0.169 0.299
Replay (Exp) + AR 0.053 0.058 0.062 0.102 0.066 0.240
Replay ($\alpha=0.1$) + AR 0.031 0.037 0.049 0.186 0.165 0.294
Replay ($\alpha=0.3$) + AR 0.035 0.042 0.050 0.157 0.136 0.282
Replay ($\alpha=0.5$) + AR 0.034 0.050 0.052 0.129 0.098 0.277
Replay ($\alpha=0.7$) + AR 0.044 0.059 0.056 0.113 0.079 0.270
Replay ($\alpha=0.9$) + AR 0.047 0.058 0.054 0.101 0.066 0.270

### v3

@@ 7 Conclusion
This paper presents work whose goal is to advance the efficiency of training and updating large language models. These improvements could help democratize LM development by enabling resource-constrained organizations to participate in research and development, while reducing the environmental footprint of model training through decreased energy consumption. However, increased accessibility of LM training also increases any risks that pertain to language models as a technology, such as the potential for malicious actors to spread disinformation.
@@ Acknowledgements
We are highly grateful to the many developers of the OpenLM library which was used extensively in this work. In particular, we’d like to thank George Smyrnis, Achal Dave, and Samir Yitzhak Gadre for helpful discussions regarding its usage. We’d also like to thank Chun-Liang Li, Pavan Kumar Anasosalu Vasu, Preetum Nakkiran, and Russ Webb for helping review earlier versions of the manuscript.
@@ B.4 Evaluation metrics
For downstream evaluations where evaluation periods do not align with model checkpoint dates, we use an exclusive assignment strategy to define in-domain (ID) performance. For each checkpoint $i$, we identify its nearest preceding evaluation – the evaluation timestep that comes before or concurrent with checkpoint $i$ and is closest in time. However, since multiple checkpoints may share the same nearest preceding evaluation, we ensure each evaluation timestep is claimed as “concurrent” by only the temporally closest training timestep.
Formally, let $t_{i}$ and $e_{j}$ denote the timestamps of checkpoint $i$ and evaluation period $j$ respectively. Checkpoint $i$ contributes $R_{i,j}$ to the ID average only if evaluation period $j$ is both (1) the nearest preceding evaluation to checkpoint $i$, and (2) checkpoint $i$ is the closest checkpoint to evaluation period $j$ among all checkpoints that have $j$ as their nearest preceding evaluation. All other pairs $R_{i,j}$ contribute to forward transfer when $t_{i}<e_{j}$, and backward transfer when $t_{i}>e_{j}$.
@@ Appendix C Training and hyperparameter details
Figure 15: Cyclic cosine learning rate schedules compared with a single cosine learning rate schedule (gray) restart the schedule to either the initial maximum (green) or gradually decaying values (blue).
@@ D.5 TiC-StackE-English
Table 8: Results for TiC-StackE-English (3B models, 440B training tokens)
@@
Cyclic Cosine 440B 0.035 0.044 0.070
Cyclic Cosine + AR 440B 0.028 0.042 0.070
Cyclic Rsqrt 440B 0.028 0.041 0.068
ScheduleFree 440B 0.029 0.044 0.071
Replay ($\alpha=1/t$) 440B 0.030 0.050 0.074
Replay ($\alpha=1/2$) 440B 0.030 0.045 0.072
Replay ($\alpha=1/t$) + AR 440B 0.030 0.052 0.075
Replay ($\alpha=1/2$) + AR 440B 0.026 0.045 0.070
LwF 440B 0.032 0.042 0.070
EWC 440B 0.030 0.043 0.069
Oracle Series 1.16T 0.018 0.037 0.074
D.6 Extended replay exploration
Table 9: Further exploration of replay variants. Results are for 3B models trained for 220B total tokens.
@@ D.6 Extended replay exploration
Replay ($\alpha=1/t$) 0.038 0.057 0.091 0.035 0.050 0.074
Replay (Exp) 0.033 0.046 0.086 0.037 0.048 0.072
Replay ($\alpha=0.1$) 0.036 0.054 0.090 0.035 0.050 0.075
Replay ($\alpha=0.3$) 0.034 0.050 0.088 0.034 0.048 0.074
Replay ($\alpha=0.5$) 0.034 0.050 0.088 0.034 0.048 0.074
Replay ($\alpha=0.7$) 0.032 0.046 0.085 0.036 0.048 0.072
Replay ($\alpha=0.9$) 0.033 0.044 0.085 0.039 0.048 0.072
Replay ($\alpha=1/t$) + AR 0.039 0.060 0.092 0.034 0.052 0.077
Replay (Exp) + AR 0.033 0.048 0.088 0.034 0.046 0.073
Replay ($\alpha=0.1$) + AR 0.038 0.056 0.091 0.034 0.049 0.076
Replay ($\alpha=0.3$) + AR 0.036 0.052 0.090 0.033 0.047 0.075
Replay ($\alpha=0.5$) + AR 0.034 0.050 0.088 0.033 0.047 0.074
Replay ($\alpha=0.7$) + AR 0.033 0.048 0.087 0.033 0.046 0.073
Replay ($\alpha=0.9$) + AR 0.033 0.046 0.086 0.035 0.047 0.072
@@
Replay ($\alpha=1/t$) 0.075 0.119 0.191 -0.009 -0.010 0.006
Replay (Exp) 0.054 0.088 0.170 0.026 0.013 0.006
Replay ($\alpha=0.1$) 0.062 0.104 0.183 -0.007 -0.008 -0.004
Replay ($\alpha=0.3$) 0.060 0.100 0.176 -0.001 -0.002 -0.001
Replay ($\alpha=0.5$) 0.055 0.090 0.170 0.010 0.002 0.002
Replay ($\alpha=0.7$) 0.050 0.086 0.169 0.018 0.009 0.005
Replay ($\alpha=0.9$) 0.042 0.076 0.160 0.028 0.016 0.007
Replay ($\alpha=1/t$) + AR 0.066 0.116 0.193 -0.019 -0.015 -0.008
Replay (Exp) + AR 0.044 0.088 0.174 0.005 0.000 0.000
Replay ($\alpha=0.1$) + AR 0.057 0.107 0.188 -0.018 -0.013 -0.006
Replay ($\alpha=0.3$) + AR 0.052 0.099 0.179 -0.012 -0.010 -0.004
Replay ($\alpha=0.5$) + AR 0.047 0.091 0.176 -0.006 -0.006 -0.002
Replay ($\alpha=0.7$) + AR 0.042 0.086 0.170 -0.001 -0.003 -0.002
Replay ($\alpha=0.9$) + AR 0.033 0.076 0.161 0.005 $0$0.001
@@
Replay ($\alpha=1/t$) 0.054 0.055 0.057 0.172 0.183 0.275
Replay (Exp) 0.070 0.068 0.063 0.069 0.087 0.233
Replay ($\alpha=0.1$) 0.042 0.045 0.051 0.163 0.171 0.268
Replay ($\alpha=0.3$) 0.056 0.059 0.057 0.130 0.139 0.250
Replay ($\alpha=0.5$) 0.058 0.062 0.060 0.097 0.110 0.237
Replay ($\alpha=0.7$) 0.066 0.073 0.068 0.087 0.101 0.232
Replay ($\alpha=0.9$) 0.069 0.072 0.064 0.066 0.083 0.225
Replay ($\alpha=1/t$) + AR 0.040 0.042 0.057 0.195 0.202 0.277
Replay (Exp) + AR 0.053 0.052 0.062 0.099 0.114 0.240
Replay ($\alpha=0.1$) + AR 0.031 0.032 0.049 0.184 0.190 0.271
Replay ($\alpha=0.3$) + AR 0.035 0.036 0.050 0.155 0.161 0.254
Replay ($\alpha=0.5$) + AR 0.035 0.038 0.052 0.127 0.138 0.246
Replay ($\alpha=0.7$) + AR 0.044 0.046 0.056 0.111 0.121 0.238
Replay ($\alpha=0.9$) + AR 0.048 0.045 0.054 0.098 0.112 0.234
@@
D.7 Deduplicating newer against older data
In this section, we explore an alternative approach to data deduplication. As mentioned in Section 3, for all other experiments, we performed deduplication only within each month. Here, we consider the effects of also deduplicating newer months against older months. Specifically, we implement this by simply retaining the state of the Bloom Filter across months, only resetting it once every 10 months, which we refer to as a "windowed" deduplication.
Table 10: Exploration of windowed deduplication (3B models, 440B training tokens). Here “Loc.” refers to the original within-month local deduplication strategy and “Wind.” refers to our 10-month windowed deduplication.
[…]
Bwd. ID Fwd. Bwd. ID Fwd. Bwd. ID Fwd. Bwd. ID Fwd.
Replay ($\alpha_{t}=1/2$) Loc. 0.015 0.029 0.073 0.017 0.027 0.055 0.027 0.065 0.158 0.020 0.032 0.189
Replay ($\alpha_{t}=1/2$) + AR Loc. 0.010 0.027 0.073 0.007 0.022 0.054 0.036 0.074 0.160 0.058 0.070 0.214
Replay ($\alpha_{t}=1/2$) Wind. 0.013 0.026 0.072 0.014 0.023 0.052 0.029 0.059 0.149 0.036 0.048 0.199
Replay ($\alpha_{t}=1/2$) + AR Wind. 0.007 0.024 0.070 0.002 0.017 0.049 0.037 0.071 0.154 0.064 0.078 0.221
Oracle Series Loc. 0.014 0.035 0.080 0.013 0.030 0.061 0.012 0.056 0.146 0.035 0.057 0.196
Method Dedup TiC-StackE-Math $\downarrow$ TiC-CD-NumPy $\downarrow$ Static Evals. $\uparrow$
Bwd. ID Fwd. Bwd. ID Fwd. Core (DCLM)
Replay ($\alpha_{t}=1/2$) Loc. $0$0.008 -0.011 0.029 0.038 0.044 50.1
Replay ($\alpha_{t}=1/2$) + AR Loc. -0.015 -0.019 -0.017 0.036 0.039 0.043 49.2
Replay ($\alpha_{t}=1/2$) Wind. 0.011 0.002 -0.001 0.085 0.088 0.086 49.8
Replay $\alpha_{t}=1/2$) + AR Wind. -0.002 -0.009 -0.008 0.078 0.078 0.078 49.5
Oracle Series Loc. -0.025 -0.028 -0.022 0.008 0.008 0.015 50.6
In Table 10, we apply the new deduplication strategies and measure performance on downstream evaluations only (as TiC-CC evaluations will be biased towards the original within-month deduplication as it was used to construct TiC-CC test sets). Overall, we observe that deduplicating across months, when applied with replay, can help improve performance on some evaluations but not all. We speculate that this is because on rarer domains, such as code, deduplicating more aggressively across months can reduce the already limited amount of relevant samples. On the other hand, for evaluations where more general data helps, such as TiC-Wiki, deduplication helps by letting the model see more diverse data given the same token budget.
D.8 Effects of model size on downstream evaluations.
Table 11: Comparing downstream evaluations for 1B and 3B models (220B training tokens). For all dynamic evaluations, we report perplexity values relative to the Oracle-2024-07 model of the same size and with log-scaling. Meanwhile, Core is an average of the accuracies of 22 downstream zero/few-shot tasks used by Li et al. (2024a), evaluated only on the final model checkpoint (without scaling by an Oracle). Bold values are within one standard deviation (estimated with 3 runs of Cyclic Cosine) of the best in each column for a given model size.
[…]
Bwd. ID Fwd. Bwd. ID Fwd. Bwd. ID Fwd. Bwd. ID Fwd.
Cyclic Cosine (std) 3B 0.033 0.045 0.085 0.039 0.048 0.072 0.041 0.073 0.156 0.057 0.072 0.219
Cyclic Cosine + AR 3B 0.033 0.048 0.087 0.035 0.047 0.074 0.032 0.072 0.159 0.081 0.096 0.228
Cyclic Rsqrt 3B 0.031 0.043 0.084 0.035 0.045 0.070 0.035 0.071 0.158 0.059 0.074 0.220
Schedule-Free 3B 0.035 0.048 0.087 0.040 0.050 0.074 0.038 0.074 0.160 0.081 0.096 0.236
Replay ($\alpha_{t}=1/t$) 3B 0.038 0.057 0.091 0.035 0.050 0.074 0.075 0.119 0.191 0.172 0.183 0.275
Replay ($\alpha_{t}=1/2$) 3B 0.032 0.047 0.086 0.034 0.047 0.072 0.055 0.090 0.170 0.097 0.110 0.237
Replay ($\alpha_{t}=1/t$) + AR 3B 0.039 0.060 0.092 0.034 0.052 0.077 0.066 0.116 0.193 0.195 0.202 0.277
Replay ($\alpha_{t}=1/2$) + AR 3B 0.034 0.050 0.088 0.033 0.047 0.074 0.047 0.091 0.176 0.127 0.138 0.246
LwF 3B 0.033 0.045 0.085 0.039 0.048 0.072 0.037 0.070 0.155 0.055 0.070 0.216
EWC 3B 0.031 0.044 0.083 0.034 0.045 0.069 0.033 0.072 0.162 0.067 0.080 0.222
Oracle Series 3B 0.014 0.035 0.080 0.013 0.030 0.061 0.012 0.056 0.146 0.035 0.057 0.196
Method Params TiC-StackE-Math $\downarrow$ TiC-CD-NumPy $\downarrow$ Static Evals. $\uparrow$
Bwd. ID Fwd. Bwd. ID Fwd. Core (DCLM)
Cyclic Cosine (std) 3B 0.037 0.024 0.015 0.070 0.075 0.070 48.5
Cyclic Cosine + AR 3B 0.011 0.006 0.003 0.055 0.056 0.062 48.5
Cyclic Rsqrt 3B 0.024 0.016 0.009 0.067 0.072 0.071 49.0
Schedule-Free 3B 0.035 0.025 0.017 0.070 0.075 0.080 48.8
Replay ($\alpha_{t}=1/t$) 3B -0.009 -0.010 -0.006 0.054 0.055 0.057 48.9
Replay ($\alpha_{t}=1/2$) 3B 0.010 0.002 0.002 0.058 0.062 0.060 49.0
Replay ($\alpha_{t}=1/t$) + AR 3B -0.019 -0.015 -0.008 0.040 0.042 0.057 49.0
Replay ($\alpha_{t}=1/2$) + AR 3B -0.006 -0.006 -0.002 0.035 0.038 0.052 49.2
LwF 3B 0.034 0.022 0.013 0.073 0.078 0.073 48.5
EWC 3B 0.016 0.010 0.006 0.056 0.060 0.067 48.9
Oracle Series 3B -0.025 -0.028 -0.022 0.008 0.008 0.015 50.6
[…]
Bwd. ID Fwd. Bwd. ID Fwd. Bwd. ID Fwd. Bwd. ID Fwd.
Cyclic Cosine (std) 1B 0.035 0.044 0.079 0.040 0.045 0.064 0.074 0.113 0.211 0.061 0.079 0.232
Cyclic Cosine + AR 1B 0.035 0.045 0.081 0.035 0.043 0.065 0.089 0.135 0.222 0.104 0.120 0.252
Cyclic Rsqrt 1B 0.034 0.044 0.079 0.037 0.044 0.063 0.084 0.126 0.219 0.082 0.099 0.245
ScheduleFree 1B 0.040 0.049 0.083 0.044 0.050 0.068 0.095 0.138 0.232 0.111 0.127 0.262
Replay ($\alpha=1/t$) 1B 0.038 0.054 0.085 0.034 0.046 0.066 0.100 0.152 0.236 0.185 0.196 0.284
Replay ($\alpha=1/2$) 1B 0.036 0.048 0.081 0.036 0.045 0.064 0.089 0.129 0.218 0.115 0.129 0.248
Replay ($\alpha=1/t$) + AR 1B 0.039 0.055 0.085 0.032 0.044 0.066 0.113 0.167 0.245 0.214 0.225 0.299
Replay ($\alpha=1/2$) + AR 1B 0.036 0.049 0.083 0.032 0.043 0.064 0.097 0.145 0.231 0.152 0.165 0.272
LwF 1B 0.036 0.045 0.080 0.041 0.046 0.065 0.076 0.119 0.213 0.065 0.083 0.237
EWC 1B 0.033 0.043 0.078 0.035 0.042 0.061 0.086 0.128 0.220 0.097 0.114 0.249
Oracle Series 1B 0.015 0.043 0.073 0.016 0.040 0.058 0.032 0.089 0.176 0.040 0.000 0.201
Method Params TiC-StackE-Math $\downarrow$ TiC-CD-NumPy $\downarrow$ Static Evals. $\uparrow$
Bwd. ID Fwd. Bwd. ID Fwd. Core (DCLM)
Cyclic Cosine 1B 0.063 0.051 0.034 0.097 0.115 0.123 45.3
Cyclic Cosine + AR 1B 0.042 0.033 0.023 0.093 0.112 0.126 45.6
Cyclic Rsqrt 1B 0.054 0.043 0.029 0.096 0.115 0.127 45.7
ScheduleFree 1B 0.067 0.055 0.041 0.112 0.132 0.142 45.6
Replay ($\alpha=1/t$) 1B 0.013 0.012 0.011 0.082 0.095 0.110 45.6
Replay ($\alpha=1/2$) 1B 0.032 0.024 0.018 0.084 0.097 0.110 45.6
Replay ($\alpha=1/t$) + AR 1B 0.008 0.010 0.011 0.090 0.103 0.116 45.6
Replay ($\alpha=1/2$) + AR 1B 0.021 0.016 0.015 0.082 0.099 0.114 45.9
LwF 1B 0.062 0.049 0.034 0.096 0.119 0.130 45.6
EWC 1B 0.041 0.033 0.023 0.107 0.123 0.128 45.6
Oracle Series 1B -0.012 -0.017 -0.016 0.016 0.025 0.037 46.7
As seen in Table 11 above, for the same token budget, changing model size can lead to continual methods being more suboptimal compared to Oracle-2024-07 on all evaluations besides TiC-Wiki. However, the overall conclusions do not majorly change (e.g., which evaluations benefit from replay versus not).